OpenStudy (anonymous):

Convert the integrals into polar coordinate form. x is from 0 to 2, y is o to sqrt(2x-x^2)?

OpenStudy (amistre64):

r=sqrt(x^2+y^2)

OpenStudy (amistre64):

and theta = tan-1(y/x)

OpenStudy (amistre64):

r = sqrt(x^2+2x-x^2) = sqrt(2x)

OpenStudy (anonymous):

what does tan-1(y/x) mean?

OpenStudy (amistre64):

arctangent function

OpenStudy (amistre64):

tan(t) = y/x t = tan-1(y/x)

OpenStudy (anonymous):

The question ask to integrate those over 5*sqrt(x^2+y^2) though

OpenStudy (anonymous):

Those were the integrals.

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=polar+plot%28sqrt%282x%29%2C%28arctan%28sqrt%282x-x%5E2%29%29%28x%29%29%29 then you prolly werent want to convert into polar coordinates then

OpenStudy (amistre64):

most likely you have a vector valued function and want to integrate it

OpenStudy (anonymous):

Question is "Evaluate the iterated integral by converting to polar coordinates."

OpenStudy (anonymous):

the function is 5*sqrt(x^2+y^2), and the integrals are 0 to 2 for x, 0 to sqrt(2x-x^2) for y, dy dx.

OpenStudy (amistre64):

sounds vaguely familiar

OpenStudy (amistre64):

\[\int_{0}^{2}\int_{0}^{2x-x^2}5\sqrt{x^2+y^2}\ dydx\]

OpenStudy (anonymous):

2x-x^2 is square-rooted

OpenStudy (amistre64):

r = sqrt(x^2+y^2 sooo) 5r \[\int_{0}^{2}\int_{0}^{\sqrt{2x-x^2}}5r\ dydx\] and know your asking to change the limits ....

OpenStudy (amistre64):

i believe r = sqrt(2x) is still valid for dr

OpenStudy (amistre64):

oy!! 5 sqrt(x^2+y^2 ) is a halfcircle centered at the origin with a radius of 5

OpenStudy (anonymous):

.....And then what? :o

OpenStudy (anonymous):

integral would be 0 to pi/2 and 0 to 5?

OpenStudy (anonymous):

But the answer doesn't have any pi, which is very weird!

OpenStudy (amistre64):

this is a cone in 3d space http://web.monroecc.edu/manila/webfiles/calcNSF/JavaCode/CalcPlot3D.htm

OpenStudy (amistre64):

helps if i can see these things right :)

OpenStudy (anonymous):

Sorry dude but I really don't get it. But thanks for trying to get me to understand it

OpenStudy (amistre64):

'sok, still kinda fuzzy on these myself

OpenStudy (amistre64):

 this is a rough idea of our domain

OpenStudy (amistre64):

and "above" this is our surface 5sqrt(x^2+y^2)

OpenStudy (anonymous):

Yup i got that part actually. i just cant figure out the integrals! :/

OpenStudy (amistre64):